A 2.01 μFμF capacitor that is initially uncharged is connected in series with a 6.51 kΩkΩ resistor and an emf source with 74.6 VV and negligible internal resistance. The circuit is completed at tt = 0.

Just after the circuit is completed, what is the rate at which electrical energy is being dissipated in the resistor?

At what value of tt is the rate at which electrical energy is being dissipated in the resistor equal to the rate at which electrical energy is being stored in the capacitor?

At the time calculated in part B, what is the rate at which electrical energy is being dissipated in the resistor?
Answer #1

R C = 2.01 MF V = 74.6V R= 6.51 kr. 4 min Le Recall q = CUCI - VRC) = v= V(i-et lec) so i = v. ERC current through capacitor.